OpenStudy (anonymous):

2 questions: 4 bananas are to be selected from a group of 6 how many ways can this be done. 2nd question is the same thing but instead of 6 its 9

OpenStudy (anonymous):

bananas all look the same to me :)

OpenStudy (anonymous):

gtfo

OpenStudy (anonymous):

Serious. You have to pick from distinguishable objects. Now let us assume the bananas have numbers paintedon them, or something.

OpenStudy (anonymous):

Becomes a simple combinations problem. do you know the formula?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

if i knew this formula iwouldnt of asked for hel

OpenStudy (anonymous):

formula schmormula pick one banana, there are 6 ways to do it pick another, there are now 5 choices counting principle says there are \(6\times 5\) ways to do this altogether, but since you cannot tell the bananas apart you have counted too many ways, since picking banana 1 and then banana 2 is the same as picking banana 2 and then banana 1 therefore you have overcounted by a factor of two and the answer is \(\frac{6\times 5}{2}=3\times 5=15\)

OpenStudy (anonymous):

@satellite73 doesn't like formulas there is something to that.

OpenStudy (anonymous):

so what if you can pick four bananas from a group of 9?

OpenStudy (anonymous):

pick one banana, there are 9 ways to do it pick another, there are now 8 choices counting principle says there are 9×8 ways to do this altogether, but since you cannot tell the bananas apart you have counted too many ways, since picking banana 1 and then banana 2 is the same as picking banana 2 and then banana The answer is \[\frac{n!}{(n-k)! \ k!}\] where n = 9 and k = 4

OpenStudy (anonymous):

i didnt get the right answer wtf

OpenStudy (anonymous):

I calculate >>> 9*8*7*6/4*3*2 4536

OpenStudy (anonymous):

\[\dbinom{9}{4}=\frac{9\times 8\times 7\times 6}{4\times 3\times 2}=3\times 7\times 6=126\]